How to caculate the return on the mutual funds?
Answer #1

Suppose you have bought 100 units of a mutual fund at NAV (Net asset value ) Rs 10/-.

So your total investment is = 100*10 = Rs 1000/-

Now if after one year, the NAV of the fund increases to Rs 12/-

Then you have a gain of Rs 2/- (12-10)

If you redeem all of your units 100 units now you will get= 100*12= Rs 1200/-

Hence your gain will be = Redemption value - Investment value

= 1200-1000 = Rs 200/-

In percentage terms your return is = gain investment * = | 200 - * 100 = 20 1000

So the return is 20%